Devon tosses a horseshoe at a stake 30 feet away. The horseshoe lands no more than 3 feet from the stake. (a) Write an absolute value inequality that represents the range of distances that the horseshoe travels. (b) Solve the inequality.
Answer by stanbon(75887)

Devon tosses a horseshoe at a stake 30 feet away. The horseshoe lands no more than 3 feet from the stake. (a) Write an absolute value inequality that represents the range of distances that the horseshoe travels. (b) Solve the inequality.
----------
Let "x" be the distance the horseshoe travels.
|x-30|<=3
-3 <= x-30 <=3
30-3 <= x <= 30+3
27 <= x <= 33
===================
Cheers,
Stan H.